OpenStudy (anonymous):

Which of the following is equivalent to 48xy2 + 24xy4 − 12x2y4?

OpenStudy (anonymous):

12xy2(4 + 2y2 − xy2) 12xy2(4 − 2y2 + xy2) 12xy2(3 + 2y2 − xy2) 12xy2(4 − 3y2 − xy2)

OpenStudy (mathmale):

A-G: How about looking for factors common to all 3 terms of the expression 48xy2 + 24xy4 − 12x2y4?

OpenStudy (mathmale):

For example, what do 48, 24 and -12 have in common?

OpenStudy (anonymous):

2

OpenStudy (mathmale):

Certainly, but what about 3? 4? 6? 12?

OpenStudy (mathmale):

Hint: 48xy2 + 24xy4 − 12x2y4 = 12( ??? )

OpenStudy (anonymous):

1

OpenStudy (anonymous):

or 3 rite

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

48xy2 + 24xy4 − 12x2y4 = 12( ??? ) asks you to decide what the right side of the equation looks like after you've factored 12 out of 48xy2 + 24xy4 − 12x2y4. What's 48/12? What's 24/12? What's -12x^2y^4? Then 48xy2 + 24xy4 − 12x2y4 = 12( ????? )

OpenStudy (anonymous):

im lost

OpenStudy (mathmale):

Here's an example of what I'm asking you to do: 48xy2 + 24xy4 − 12x2y4 =\[12(\frac{ 48xy ^{2} }{12 } + \frac{ 24xy ^{4} }{ 12 } + ... )\]

OpenStudy (mathmale):

This simplifies to \[12(4xy ^{2}+2xy ^{4}-1x ^{2}y ^{4}).\]

OpenStudy (mathmale):

Please, AG, take a look at this and try to explain in your own words what I've done here and why. Then we'll move forward.